OpenStudy (anonymous):

Evaluate 2x-4y for x = 2 and y = 4.

OpenStudy (shamim):

2x-4y=2*2-4*4=?

OpenStudy (shamim):

But u should not give answer first!

OpenStudy (anonymous):

The only possible answers are -64, 1/8 and 1/2. And yes i agree. Don't give answers first when I'm on here for help.

OpenStudy (anonymous):

Okay I'm not on here to get made fun of. I'm dyslexic and sometimes need help. Is that so much to ask?

OpenStudy (anonymous):

And I'm not a guy.

OpenStudy (anonymous):

I apologize.

OpenStudy (anonymous):

Anyway I wish you a happy learning. I am sorry to have generalized

OpenStudy (anonymous):

Thanks for making me feel bad about myself. Have a nice day! And no. I won't like the comments because you did nothing.

OpenStudy (usukidoll):

@jordyn1212 and @shamim are right. Robert this is for you because I have seen so many rule breaking on OpenStudy that it's not funny. From the OpenStudy Code of Conduct OpenStudy is a global learning community - built for you and by you. We ask all members to abide by a few core principles, an OpenStudiers’ pledge if you will. As an OpenStudier, I will: Be Nice I will stay positive, be friendly, and not mean .... that's not all Don't be offensive, inappropriate, or creepy! So avoid: Insulting, berating, or otherwise discouraging others

OpenStudy (anonymous):

I just come on here to get help..

OpenStudy (anonymous):

Don't swear on my stuff. Please and thank you.

OpenStudy (anonymous):

And stop being so mean!

OpenStudy (anonymous):

No you will not help me with my math anymore because I'm going to block you. You did not help what so ever and made me feel bad. Math has always been a struggle for me and I was just asking for a second opinion so I could make sure my answer was right.

rvc (rvc):

good job @UsukiDoll :)